A fair coin tossed 4 times. What is te probability all tosses will be heads?
Found 2 solutions by checkley79, Lucky Duck:
Answer by checkley79(3341)

1/2*1/2*1/2*1/2=1/8=.0625=6.25% OF THE TIME.

Answer by Lucky Duck(4)

.5 * .5 * .5 * .5 = 0.0625 = 6.25% of the time.